OpenStudy (anonymous):

Find the length of the missing side. If necessary, round to the nearest tenth.

OpenStudy (anonymous):

what side?

OpenStudy (anonymous):



OpenStudy (anonymous):

@Ziegenbein

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

16^2+22^2=?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

point what because it says to round to the nearest 10th

OpenStudy (anonymous):



OpenStudy (anonymous):

27.2 right

OpenStudy (anonymous):

thats the answer? ^

OpenStudy (anonymous):

18.5^2+17.5^2=?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

648.5

OpenStudy (anonymous):

so for the answer I would write down ''27'' or ''27.2'' ?

OpenStudy (anonymous):

square root it

OpenStudy (anonymous):

27.2

OpenStudy (anonymous):

ok thats the final answer ?

OpenStudy (anonymous):

yes for the first 1

OpenStudy (anonymous):

what about the second?

OpenStudy (anonymous):

18.5^2+17.5^2=?

OpenStudy (anonymous):

?=648.5

OpenStudy (anonymous):

then you square root it

OpenStudy (anonymous):

25

OpenStudy (anonymous):

and get 25.46566316

OpenStudy (anonymous):

rounded you get 25.5

OpenStudy (anonymous):

is your final answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

25.5 would be my final answer?

OpenStudy (anonymous):

yes for the second 1

OpenStudy (anonymous):

tysm

OpenStudy (anonymous):

ily

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

now you can close the question